2. (a) Using data commonly found on the Internet about the properties of water, calculate the normal boiling point of water using this equation: lnK = - (deltaH/R)(1/T) + C
(b) What are some possible reasons that your answer may differ from the accepted value of
100 °C?
Answer #1

Following is the - complete; Answer -&- Explanation: for the given: Question: in....typed format...

\RightarrowAnswer:

  1. part - (a):  Normal boiling point of water: Tbp = 377.6 K = 104.6 oC   
  2. part - (b):  As explained below: under Step - 3, in the Explanation...

\RightarrowExplanation:

Following is the complete Explanation: for the above: Answer.

  • Given:
  1. Given:  clausius-clapeyron equation:     ln K = - ( \Delta Hvap / R  ) + C
  2. We know: vapor pressure of pure water : P1 = 23.8 mmHg = 23.8 Torr, at T1  = 298 K ( Kelvin )
  3. We know: enthalpy of vaporization of pure water:   \DeltaHvap = 40.7 kJ /mol ( approx. )
  4. We know: at normal boiling point of Water: vapor pressure of water: P2 = 760.0 Torr  ( i.e. 1.0 atm )
  5. Let's assume: normal boiling point of water: T2 = ( unknown )
  • ​​​​​​​Step - 1:

Using the Calusius-Clapeyron Equation: we can derive the following Equation:

\Rightarrowln ( P1 / P2 ) = -   \DeltaHvap / R x (  1/ T1 - 1/T2  ) -------------------------------Equation - 1

\Rightarrow Where:

  1. P1 =   23.8 Torr,
  2. P2 = 760.0 Torr ( i.e. 1.0 atm. )
  3.   \DeltaHvap =   40.7 kJ /mol ( approx. ) = 40,700 J / mol
  4. Ideal Gas constant: R = 62.364 L. Torr/ ( mol. K ) = 8.314 J / (mol. K )
  5. T1 = 298.0 K  ( Kelvin )
  6. Normal boiling point of Water:   T2 = ( unknown )
  • ​​​​​​​Step - 2:

​​​​​​​Now, plugging in values: in : Equation - 1, we would get, the following:

\Rightarrow(  1/ T1 - 1/T2  ) = 0.0007074  

\Rightarrow   (  1/ 298   - 1/T2  ) = 0.0007074  

\Rightarrow   1/T2 =   0.002648

\Rightarrow T2 = 377.6 K = ( 377 - 273 ) =  104.6 oC

\Rightarrow Therefore: Calculated normal boiling point of water : Tbp = 104.6 oC   

  • Step - 3:

​​​​​​​\Rightarrow We know: accepted value of boiling point of water = 100.0 oC   

\RightarrowHere, the calculated value of the boiling point of water, is slightly more than the accepted value, because: the values used above: depends on the value of enthalpy of vaporization, and the value of vapor pressure at 298 K , which may have some errors. Due to the above said errors, the calculated value of the normal boiling point of Water, may vary from : the accepted value: of normal boiling point of Water...

​​​​​​​